Let us pray to the Lord, saying, Lord, have mercy. For wisdom and a sense of justice to all the world’s leaders, Let us pray to the Lord For a love and respect of this planet, that all may use its abundance wisely and in accordance with your

wishes, Let us pray to the Lord For health and healing to those who are sick or dispairing, Let us pray to the Lord For the lonely and the friendless, that they may find acceptance, Let us pray to the Lord For this broken world, that it may regain a sense of community and mutual love and acceptance, Let us

pray to the Lord

Especially in this cold weather, may all those who are poor, needy, or homeless find comfort and succor, Let us pray to the Lord

We thank you for all of our creature comforts; may we always remember those who are less fortunate than us, Bless we the Lord

We thank you for the seasons; may we always see the wonder of your creation, Bless we the Lord For all of the prayers and concerns of this gathering, whether spoken silently or alound…